# A Geometry-Grounded Data Perimeter in Azure

### v2

@@ Abstract
While data perimeter is ubiquitous in cybersecurity speak, it rarely defines how boundary points are arranged. In this paper we show how Azure’s blast radius ultrametric provides the distance, and how solving the Traveling Salesman Problem in this ultrametric space provides the ordering, yielding a true geometric contour: an actionable perimeter measure for SPN data plane operations de-risking.
@@ 2.1.2 Ultrametrics and Hierarchical Models
Our use of ultrametrics follows this tradition, but we leverage their properties for efficient tour construction, not just visualization or embedding.
@@ 2.3 Blast Radius Ultrametric
The Blast radius was introduced to address the challenge of measuring lateral motion, data leakage and data forgery of individual SPNs in Azure. It is fully impplemented in Azure Silhouette[8].
@@
The blast radius of an SPN is $B=\max_{i,j}d(x_{i},x_{j})$. It ranges from 0.0 (no permissions) to 1.0 (Tenant-wide permissions).
@@ 4.1 Formulation
So, the data perimeter is not only geometrically grounded but also computationally tractable in our setting. The polynomial complexity allows for practical scaling even with large sets of data actions or SPNs. It is the sum of distances in the tour, with wraparound from the last data point to the first:
$P=\sum_{i=1}^{n}d(x_{i},x_{i+1}),\quad\text{with }x_{n+1}=x_{1}$
Note that:
@@ 4.1.1 Ultrametric TSP pseudo-code
    Find $a_{\text{start}}\in A$ such that
    $a_{\text{start}}=\arg\max_{a\in A}\max_{b\in A}d(a,b)$
@@ 4.1.2 Limitation
When hierarchical families are defined, Silhouette computes the Blast radius as the pointwise infimum in all hierarchies to yield a tighter, smaller Blast Radius than the native one. In doing so, it must be noted that the pointwise infimum is not ultrametric, hence the data Perimeter should be treaded with care in this situation. Falling back to the native tenant hierarchy is highly recommended.
@@
5 Mean distance
The average distance between any two data actions is the metric that springs to mind for complementing the Blast radius. It captures global dispersion but ignores local structure and chaining. In this section, we explain why this doesn’t make it a good secondary key candidate for sorting SPNs.
@@ 5.1 Comparison with Data Perimeter
    For each SPN we compute a spread_ratio defined as $\frac{\text{P}}{n\cdot\mu_{\mathrm{mean}}}$, where $n$ is the data actions count held by this SPN. The reason why we divide $P$ by $n$ is to make it comparable with $\mu_{\mathrm{mean}}$.
@@ 5.1.1 Discussion
When many SPNs share the same blast radius, the data perimeter provides a fine-grained, geometry-aware tiebreaker. It surfaces principals whose permission sets exhibit branching or elongated structures (profiles that pose higher risk in practice) whereas the mean alone would treat them identically to more compact SPNs.

### v3

@@ Abstract
While Data Perimeter is ubiquitous in cybersecurity speak, it rarely defines how boundary points are arranged. In this paper we show how Azure’s blast radius ultrametric provides the distance, and how solving the Traveling Salesman Problem in this ultrametric space provides the ordering, yielding a true geometric contour: an actionable perimeter measure for Service Principals (SPNs) data plane operations de-escalation.
@@ 2.1.2 Ultrametrics and Hierarchical Models
Our use of ultrametrics follows this tradition, but we leverage their properties for efficient tour construction (not just visualization).
@@ 2.3 Blast Radius Ultrametric
The Blast radius was introduced to address the challenge of measuring lateral motion, data leakage and data forgery of individual SPNs in Azure. It is fully implemented in Azure Silhouette[8].
@@
The blast radius of an SPN is the diameter of $d()$: $\max_{i,j}d(x_{i},x_{j})$. It ranges from 0.0 (no permissions) to 1.0 (Tenant-wide permissions).
@@ 4.1 Formulation
As will be shown in the ultrametric TSP pseudo-code below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
@@ 4.1.1 Ultrametric TSP pseudo-code
    Pick $a_{\text{start}}\in A$ at random
@@ 4.1.3 Known Limitations
When hierarchical families are defined, Silhouette computes the Blast radius as the pointwise infimum in all hierarchies to yield a tighter, smaller Blast Radius than the native one. In doing so, it must be noted that the pointwise infimum is not ultrametric, hence the data Perimeter should be treated with care in this situation. Falling back to the native tenant hierarchy is highly recommended.
@@
5 Mean distance ($\mu_{\mathrm{mean}}$)
The average distance between any two data actions is the metric that springs to mind for complementing the Blast radius. Unfortunately, if it captures global dispersion, it ignores local structure and chaining. In this section, we explain why this doesn’t make it a good secondary key candidate for sorting SPNs.
@@ 5.1 Comparison with Data Perimeter
    For each SPN we compute a spread_ratio defined as $\frac{\text{P}}{n\cdot\mu_{\mathrm{mean}}}$, where $n$ is the data actions count held by this SPN. The perimeter $P$ is normalized by the action count $n$ to maintain dimensional consistency with $\mu_{\mathrm{mean}}$.
@@ 5.1.1 Discussion
When many SPNs share the same blast radius, the Data Perimeter provides a fine-grained, geometry-aware tiebreaker. It surfaces principals whose permission sets exhibit branching or elongated ultrametric structures whereas the mean alone would treat them identically to more compact SPNs.
@@ 6 Conclusion
Appendix: Minimal Data Perimeter
In this work, we have argued that the Data Perimeter in Azure security models behaves not just as a metaphor, but as a measurable geometric quantity. To further give substance to this geometric view, we identify the shapes corresponding to the smallest Data Perimeter.
Ultracycle and Minimal Perimeter
Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\delta>0$.
In any ultrametric space, ultracycles uniquely minimize the TSP tour length among all configurations with the same number of points.
Proof
Let $(X,d)$ be an ultrametric space and consider $n$ points $\{x_{1},x_{2},\ldots,x_{n}\}\subset X$.
Case 1: Ultracycle Configuration
Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all $i\neq j$.
For any TSP tour $\pi$ visiting all points exactly once, the tour length is:
$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\delta=n\delta$ (1)
where $x_{\pi(n+1)}:=x_{\pi(1)}$.
Since all edges have the same weight $\delta$, every possible tour has identical length $n\delta$.
Case 2: Non-Ultracycle Configuration
Now suppose the configuration is not an ultracycle. Then there exist distinct indices $i,j,k,\ell$ such that:
$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$ (2)
Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$ and $d_{\max}=\max_{i\neq j}d(x_{i},x_{j})$.
By assumption, $d_{\min}<d_{\max}$.
Lemma: in any ultrametric space, if $d(a,b)<d(a,c)$, then $d(b,c)=d(a,c)$.
Proof of Lemma: By the strong triangle inequality: $d(a,c)\leq\max\{d(a,b),d(b,c)\}$.
So we must have $d(a,c)\leq d(b,c)$.
Similarly, $d(b,c)\leq\max\{d(b,a),d(a,c)\}$. But $d(a,b)<d(a,c)$, so $d(b,c)\leq d(a,c)$.
Therefore, $d(b,c)=d(a,c)$. This proves the Lemma.
Now, let $S_{\min}=\{(i,j):d(x_{i},x_{j})=d_{\min}\}$ be the set of point pairs with minimum distance.
Claim: Any TSP tour must include at least one edge with distance $>d_{\min}$.
Proof of Claim: Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$. This would mean we can connect all $n$ points using only edges from $S_{\min}$, forming a Hamiltonian cycle.
However, by the ultrametric property (Lemma), if $d(x_{i},x_{j})=d_{\min}$ and $d(x_{j},x_{k})=d_{\min}$, then either:
  • $d(x_{i},x_{k})=d_{\min}$ (all three points are equidistant), or
  • $d(x_{i},x_{k})>d_{\min}$ (the points form a non-equidistant triple)
If all points were connected by minimum-distance edges, then by transitivity of the equidistance relation in ultrametric spaces, all pairwise distances would equal $d_{\min}$, contradicting our assumption that the configuration is not an ultracycle.
Therefore, any Hamiltonian cycle must include at least one edge with distance $>d_{\min}$.
Completing the Proof:
Let $P_{\text{non-ultra}}$ be the length of an optimal TSP tour in the non-ultracycle configuration.
Since the tour must include at least one edge of length $>d_{\min}$, and the remaining $(n-1)$ edges have length $\geq d_{\min}$:
$P_{\text{non-ultra}}>d_{\min}+(n-1)\cdot d_{\min}=n\cdot d_{\min}$ (3)
For an ultracycle with the same minimum distance $d_{\min}$ (i.e., $\delta=d_{\min}$):
$P_{\text{ultra}}=n\cdot d_{\min}$ (4)
Therefore: $P_{\text{non-ultra}}>P_{\text{ultra}}$.
This proves that ultracycles uniquely minimize the TSP tour length in ultrametric spaces.
Security Implications
In Azure’s permission model, a Principal with $n$ data actions all scoped to the same resource (forming an ultracycle at depth $d$ in the hierarchy) achieves the minimal Data Perimeter, representing optimal privilege containment.
This geometric result aligns with the theoretical foundation for the least-privilege principle: permissions should be clustered as tightly as possible in the resource hierarchy to minimize potential lateral movement exposure.